Could I get help with two questions surrounding joules? A battery-powered train travels in a complete circle of radius 1.82 m. The coefficient of friction between the train and the track is 0.12, and the train's weight is  How much work does the train do against friction with the track?   0 J 370 J 89.2 J 1800 J An object that weighs 10 N is released at a height of 5 m above the ground. After it has fallen 1 meter, what is its total mechanical energy with respect to the ground?   15 J 59 J 50 J 32 J
Answer #1

Here

radius , r = 1.82 m

u = 0.12

W = 65 N

work done by friction = distance * force

work done by friction = 65 * 0.12 * 2pi * 1.82

work done by friction = 89.2 J

-------------------------

W = 10 N

h = 5 m

as the total mechanical energy is conserved

after falling 1 m

total mechanical energy = W * h

total mechanical energy = 10 * 5

total mechanical energy = 50 J